# No Free Prune: Information-Theoretic Barriers to Pruning at Initialization

### v1

@@ 5.3 Finite Setting
As in Bubeck & Sellke (2023), we begin in setting where $T$ (and thus $|\mathcal{F}|$) is finite. For each $\mathbf{m}\in\{0,1\}^{p}$, define $\mathcal{W}_{\mathbf{m}}=\{\mathbf{w}\mid\exists f^{i}=(\mathbf{m}^{i},\mathbf{w}^{i})\in\mathcal{F}\text{ s.t. }\mathbf{m}^{i}=\mathbf{m}\}$, the possible weightings of the network once a mask is fixed, and additionally let $N_{\mathbf{m}}=|\mathcal{W}_{\mathbf{m}}|$. In this section, we denote $g(x)=\mathbb{E}[y\mid x]$ as the target function to learn and $z_{i}=y_{i}-g(x_{i})$ as the noise. We then obtain the following result:
Theorem 3.
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f^{W}(x_{i}))^{2}\leq\sigma^{2}-\epsilon\right)\leq(2k+2)e^{-\frac{n\epsilon^{2}}{8^{3}k}}\\ +\max\left(\frac{2^{7}a_{1}^{2}C_{0}I(\{X_{t}\}_{t\in T};W)}{\epsilon^{2}},2e^{-\epsilon^{2}/(2^{7}a_{1}^{2}C_{0})}\right),$
[…]
This first theorem illustrates that the quantity of interest that dictates the size of the model class is now $p_{\text{eff}}$, which we refer to as the effective parameter count. We sketch of the proof Theorem 3 below. We begin with the following lemma:
Lemma 2.
@@
In essence, for the chosen function $f^{W}$ to fit below the noise level, it must have large correlation with the noise $z_{i}$. It then remains to bound the probability of correlating with the noise. Define $X_{t}=\frac{1}{n}\sum_{i=1}^{n}(f^{t}(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f^{t}])z_{i}$. That this quantity for a fixed function $f^{t}$ is subgaussian is a consequence of isoperimetry - one observes from Definition 1 that an appropriate scaling of $(f^{t}(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f^{t}])$ must be $O(1)$-subgaussian. The new element of our proof is to use Lemma 1 to control this term rather than using standard tail bounds coupled with a union bound over all of $\mathcal{F}$. Explicitly, we apply the lemma taking $X_{t}$ to be this correlation term, which must be large for the function to fit below the noise level, and hence its upper bound must be large, rather than taking it to be a generalization error (which one usually hopes to be small).
@@
Lemma 3.
$I(W;\{X_{t}\}_{t\in T})\leq I(W;\mathcal{D})=I(f^{W};\mathcal{D})\leq I(\mathbf{m}^{W};\mathcal{D})+\mathbb{E}\left[\log_{2}N_{\mathbf{m}^{W}}\right]$
Note that the final expression is exactly our effective parameter count $p_{\text{eff}}$. Combining these two lemmas and Lemma 1 obtains Theorem 3 (see Appendix B).
@@ 5.4 Main Result
[…]
@@
[…]
The proof of the above follows from Theorem 3, coupled with a discretization argument and some careful modifications of the learning algorithm $W$. One should regard $\delta$ and $\epsilon$ as small, fixed constants independent of $n$ and $d$, at which point the theorem suggests that in order $f^{W}$ to be smooth, one must have $p_{\text{eff}}=\Omega\left(nd\right)$, recovering a result analogous to Bubeck & Sellke (2023). We remark that under standard training procedures, the above holds conditional on the specific weight initialization of the network, i.e. with $I(\mathbf{m}^{W};\mathcal{D})$ replaced with $I(\mathbf{m}^{W};\mathcal{D}\mid\text{init}=\text{init}_{0})$, since the initialization is sampled independently before training.
@@ 5.5 Discussion
Brute Force Search. Why cannot we brute force search over exponentially many subnetworks, testing each on the dataset, and claiming to have found the lottery ticket without training a dense model by the end of this process? The answer to this is that such a search creates high mutual information between the mask and the data, since it involves testing every masked network on the data and choosing the best performing mask. Thus the mutual information based bounds we present in this work hold, because the effective parameter count of the mask derived from this brute force search would be high. Of course, training the full model is a faster way to find a lottery ticket than a brute force search, since training neural networks takes subexponential time Livni et al. (2014). Algorithms attempting to learn the mask at initialization Sreenivasan et al. (2022); Ramanujan et al. (2020), before learning weights, are interesting in that this naturally increases $I(\mathbf{m}^{W};\mathcal{D})$ early on in training, although their results are not demonstrated at ImageNet scale, where lots of lottery ticket type results are known to break down Gale et al. (2019); Frankle et al. (2019).
@@
Necessary conditions. Our results present necessary but not sufficient conditions for smooth interpolation with a parameterized function class. Bombari et al. (2023) examines whether the overparameterization condition $p=\Omega(nd)$ is sufficient to guarantee smooth interpolation in various neural network regimes, finding that it is not for a random feature network, but is for a network in the NTK regime of neural network training Jacot et al. (2018). [^4] Bombari et al. (2023) do not include regularization in their random feature model, whereas Simon et al. (2023) assume optimal regularization, explaining the seemingly different conclusions about overparameterized RF regression models.
@@
Subtleties Around Effective Parameters. While our idea of effective parameters suggest that one can trade off $I(\mathbf{m}^{W};\mathcal{D})$ and parameter count, this does not imply that a learning algorithm that artificially inflates this mutual information, for instance by encoding a discretization of the data within the mask, will do well as a pruning algorithm. Our modified Law of Robustness presents a lower bound on Lipschitz constant, so that the actual Lipschitz constant could easily be higher than the lower bound given. Additionally, one can view the choice of a mask $\mathbf{m}^{W}$ followed by a choice of weights $\mathbf{w}^{W}$ constrained by this mask as having access to a collection of function classes $\{\mathcal{H}_{t}\}_{t\in T}$, where one first makes a choice $W$ (possibly depending on the data $\mathcal{D}$) of function class $\mathcal{H}_{W}$ and then a choice of function within the class. In such cases, this idea of effective parameterization accounting for both the mutual information of the choice of function class and the parameter count within the chosen class extends beyond the Law of Robustness to other settings (see Theorems 5, 6 in Appendix D), showing that the principle we introduce of trading off parameters and information is a general one.
@@ 6 Experiments
Because computing mutual information requires estimating high dimensional distributions, it is only exactly feasible in very small models Kraskov et al. (2004); Gao et al. (2018), which is precisely the setting where pruning to high sparsities leads to layer collapse Tanaka et al. (2020), where all the weights in one layer vanish. We perform such experiments for a tiny neural network in Appendix C.2. Here, we instead track $I(\mathbf{m^{W}};\mathcal{D})$, using two different proxies. Our experiments are meant to illustrate that information gained during training can affect model capacity, rather than serve as exhaustive examinations of large modern architectures.
@@ 6.1 Memorization capacity
Figure 1: Memorization capacity (train accuracy on noisy data) against sparsity level for different pruning methods. Staying higher on plots is better. Vertical gap between IMP/Magnitude-after pruning reflect additional memorization capacity on this dataset due to mutual information between mask and data. Two-hidden layer MLP on Gaussian data (a) and noisy FashionMNIST (b)
The first proxy for effective parameters is the memorization capacity of the network to fit noisy labels. We draw inspiration from a long line of work in statistical physics on the random perceptron (memorizing random labels with a linear function), and related settings Cover (1965); Gardner (1988); Talagrand (2000); Shcherbina & Tirozzi (2003); Stojnic (2013); Sur & Candès (2019); Candes & Sur (2020); El Alaoui & Sellke (2022); Montanari et al. (2024). For example Cover (1965) shows that $(2-\varepsilon)d$ randomly labelled points on the sphere are memorizable by a $d$-dimensional linear model, but $(2+\varepsilon)d$ are not. Extensions to deep networks are poorly understood, but one expects they too exhibit sharp phase transitions in the high-dimensional regime. That is, there may be a sharp threshold at which the network can no longer memorize all of the datapoints, and this threshold should be proportional to the number of effective parameters. This motivates using the ability of a network to memorize as a proxy for its effective parameter count.
We perform experiments to measure the memorizing capacity of a pruned network at various sparsity levels under differing pruning algorithms. More specifically, we consider a two-hidden layer network with ReLU activation, on a train set of points (Gaussian data in Figure 0(a) and FashionMNIST in Figure 0(b)) with noise (random Boolean labels in Figure 0(a), and corrupted images with true labels in Figure 0(b)), and compute the fraction of the training set the sparse network can memorize at each sparsity level.
One sees that for all datasets here IMP is more expressive and can memorize more points with the same number of unmasked parameters as the other subnetworks derived from different pruning algorithms. This gap in memorization capacity is due to mutual information with the data, providing evidence that the effective parameter count our theorems reason can indeed correspond to capacity of neural networks in practice. We emphasize that the y-axis is not test accuracy, but memorization capacity (train accuracy on highly noisy data): the fact that our plots give analogous behavior to plots of noiseless test-accuracy against sparsity in, e.g., Frankle et al. (2020b), is encouraging because the shape of such plots is precisely what we sought to explain. Experimental details are in Appendix C.1.
@@ 6.2 Correlation with noise
Figure 2: Ability to correlate with dataset noise over training as a proxy for network capacity. 5-layer ReLU MLP in a student-teacher task, with $\sigma^{2}=1$ noisy labels; (a): correlation with noise during IMP; training increases correlation with noise (hence $p_{\text{eff}}$), pruning then reduces this, before repeating; (b): Sweep over learning rates; (c): Sweep over amount of noise injected into gradients. This illustrates that our quantity of interest, $I(f^{W};\mathcal{D})$, and thus $p_{\text{eff}}$, is increasing due to data contained in gradients.
Our second proxy for mutual information is more direct: the correlation of the network function with the noise in labels over the course of training. By Lemma 1, we have that this correlation is a lower bound for the quantity of interest $I(f^{W};\mathcal{D})$. Overall, we see that as the network reaches low train error, it necessarily fits some of the noise, and so correlation with noise increases. We measure this quantity over training in a student-teacher task.
Figure 1(b) illustrates how increasing learning rate increases the rate at which MI is acquired, suggesting that gradients carry this informational quantity. To verify this, we can corrupt the gradients with noise, finding in Figure 1(c) that this correlational proxy for MI vanishes when we do so. Finally, Figure 1(a) tracks this proxy for MI during IMP, finding that the MI increases during iterations of pruning, decreases after pruning, and breaks down at late time. In context of our theory where $I(\mathbf{m^{W}};\mathcal{D})$ contributes to $p_{\text{eff}}$, this shows the role of iterative, rather than one-shot pruning, may be to increase the effective parameter count via gradient updates, so it can be pruned back down again. Hence the iterative nature of pruning allows the trade-off between parameters and information to be made many times.
7 Conclusion
Our theory and experiments illustrate how mutual information can be traded off with parameters so that subnetworks derived from pruning at initialization have a lower effective parameter count than those derived from pruning after training. Thus, there may be information-theoretic barriers to finding extremely sparse, trainable subnetworks without training the full neural network on data. This explains why widespread empirical efforts to prune at initialization have run into difficulties, and introduces a new information-parameter trade-off that may be of independent interest.
8 Acknowledgements
The authors thank Mansheej Paul, Gintare Karolina Dziugaite and Sitan Chen for helpful discussions.
Appendix A Appendix
Appendix B Proof of Theorem 4
@@
B.1 Overview of Finite setting, Theorem 3
B.1.1 Proof of Lemma 2
Lemma 4.
@@ B.1.1 Proof of Lemma 2
$A_{f}=\left\{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(x_{i}))^{2}\leq\sigma^{2}-\epsilon\right\}$
$B=\left\{\frac{1}{n}\sum_{i=1}^{n}z_{i}^{2}\geq\sigma^{2}-\frac{\epsilon}{6}\right\}$
$C=\left\{\frac{1}{n}\sum_{i=1}^{n}z_{i}g(x_{i})\geq-\frac{\epsilon}{6}\right\}$
$D_{f}=\left\{\frac{1}{n}\sum_{i=1}^{n}f(x_{i})z_{i}\geq\frac{\epsilon}{4}\right\}.$
@@
Follows from Lemma 2.1 of Bubeck & Sellke (2023). ∎
Lemma 5.
@@
Follows from Theorem 2 of Bubeck & Sellke (2023). ∎
@@
In particular $A_{f^{W}}\implies(B^{c}\cup C^{c})\cup E_{f^{W}}\cup F$, and using the bounds in Lemmas 4 and 5 now yields Lemma 2.
B.1.2 Proof of Lemma 1
Lemma 1 follows from Theorem 3 of Xu & Raginsky (2017). After translating notation, we have that $\mathbb{P}(|X_{w}|>\alpha)\leq\delta$ whenever
@@ B.1.2 Proof of Lemma 1
upon which rearranging yields exactly Lemma 1.555The explicit constant $a_{1}$ here is not the same as in Xu & Raginsky (2017) as a result of different definitions of “subgaussian” - in Xu & Raginsky (2017), the authors define $\sigma$-subgaussian to mean a random variable $U$ satisfies $\log\mathbb{E}[e^{\lambda(U-\mathbb{E}U)}]\leq\lambda^{2}\sigma^{2}/2$. The assumed condition $\mathbb{P}(|X|\geq t)\leq 2e^{-t^{2}/C}$, implies $\log\mathbb{E}[e^{\lambda X}]\leq\lambda^{2}(9C)/(2)$ (exercises 3.1 d and e of van Handel (2016)) so $a_{1}^{2}=72$.
B.1.3 Isoperimetry, Subgaussianity, and Crude Bounds
Having established Lemma 2, to show Theorem 3 it remains to upper bound $\mathbb{P}[E_{f^{W}}]$. We first show that $X_{t}=\frac{1}{n}\sum_{i=1}^{n}(f^{t}(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f^{t}])z_{i}$ is subgaussian. As in Bubeck & Sellke (2023), the isoperimetry assumption implies that $\sqrt{\frac{d}{c}}\frac{f(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f]}{L}$ is $2$-subgaussian. Since $|z_{i}|\leq 2$, it follows that $\sqrt{\frac{d}{c}}\frac{(f(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f])z_{i}}{L}$ is $8$-subgaussian, by Proposition 1.2 of Bubeck & Sellke (2023). Hence each $X_{t}$ is $\left(18\cdot 8\frac{cL^{2}}{dn}\right)$-subgaussian. For convenience we thus define $C_{0}=\frac{144cL^{2}}{dn}$.

### v2

@@ 5.3 Finite Setting
As in (Bubeck & Sellke, 2023), we begin in setting where $T$ (and thus $|\mathcal{F}|$) is finite. For each $\mathbf{m}\in\{0,1\}^{p}$, define $\mathcal{W}_{\mathbf{m}}=\{\mathbf{w}\mid\exists f^{i}=(\mathbf{m}^{i},\mathbf{w}^{i})\in\mathcal{F}\text{ s.t. }\mathbf{m}^{i}=\mathbf{m}\}$, the possible weightings of the network once a mask is fixed, and additionally let $N_{\mathbf{m}}=|\mathcal{W}_{\mathbf{m}}|$. In this section, we denote $g(x)=\mathbb{E}[y\mid x]$ as the target function to learn and $z_{i}=y_{i}-g(x_{i})$ as the noise. We then obtain the following result:
Theorem 5.3.
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f^{W}(x_{i}))^{2}\leq\sigma^{2}-\epsilon\right)\leq(2k+2)e^{-\frac{n\epsilon^{2}}{8^{3}k}}\\ +\max\left(\frac{2^{7}a_{1}^{2}C_{0}p_{\text{eff}}}{\epsilon^{2}},2e^{-\epsilon^{2}/(2^{7}a_{1}^{2}C_{0})}\right),$
where $p_{\text{eff}}=I(\mathbf{m}^{W};\mathcal{D})+\mathbb{E}\left[\log_{2}N_{\mathbf{m}^{W}}\right]$ and $C_{0}=\frac{144cL^{2}}{nd}$.
This first theorem illustrates that the quantity of interest that dictates the size of the model class is now $p_{\text{eff}}$, which we refer to as the effective parameter count. We sketch of the proof Theorem 5.3 below. We begin with the following lemma:
Lemma 5.4.
@@
In essence, for the chosen function $f^{W}$ to fit below the noise level, it must have large correlation with the noise $z_{i}$. It then remains to bound the probability of correlating with the noise. Define $X_{t}=\frac{1}{n}\sum_{i=1}^{n}(f^{t}(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f^{t}])z_{i}$. That this quantity for a fixed function $f^{t}$ is subgaussian is a consequence of isoperimetry - one observes from Definition 5.1 that an appropriate scaling of $(f^{t}(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f^{t}])$ must be $O(1)$-subgaussian. The new element of our proof is to use Lemma 5.2 to control this term rather than using standard tail bounds coupled with a union bound over all of $\mathcal{F}$. Explicitly, we apply the lemma taking $X_{t}$ to be this correlation term, which must be large for the function to fit below the noise level, and hence its upper bound must be large, rather than taking it to be a generalization error (which one usually hopes to be small).
@@
Lemma 5.5.
$I(W;\{X_{t}\}_{t\in T})\leq I(W;\mathcal{D})\\ =I(f^{W};\mathcal{D})\leq I(\mathbf{m}^{W};\mathcal{D})+\mathbb{E}\left[\log_{2}N_{\mathbf{m}^{W}}\right]$ (5)
Note that the final expression is exactly our effective parameter count $p_{\text{eff}}$. Combining these two lemmas and Lemma 5.2 obtains Theorem 5.3 (see Appendix A).
@@ 5.4 Main Result
    $\frac{1}{n}\sum_{i=1}^{n}(f^{W}(x_{i})-y_{i})^{2}\leq\sigma^{2}-\epsilon\\ \implies\mathrm{Lip}(f)\geq\frac{\epsilon}{96a_{1}\sqrt{2c}}\sqrt{\frac{nd\delta}{p_{\text{eff}}+\frac{\delta}{2}\log(4/\delta)}}$
@@
    $p_{\text{eff}}=I(\mathbf{m}^{W};\mathcal{D})+\mathbb{E}[\|\mathbf{m}\|_{1}]\log_{2}(1+60WJ\epsilon^{-1}).$
The proof of the above follows from Theorem 5.3, coupled with a discretization argument and some careful modifications of the learning algorithm $W$. One should regard $\delta$ and $\epsilon$ as small, fixed constants independent of $n$ and $d$, at which point the theorem suggests that in order $f^{W}$ to be smooth, one must have $p_{\text{eff}}=\Omega\left(nd\right)$, recovering a result analogous to (Bubeck & Sellke, 2023). We remark that under standard training procedures, the above holds conditional on the specific weight initialization of the network, i.e. with $I(\mathbf{m}^{W};\mathcal{D})$ replaced with $I(\mathbf{m}^{W};\mathcal{D}\mid\text{init}=\text{init}_{0})$, since the initialization is sampled independently before training.
@@ 5.5 Discussion
Brute Force Search. Why cannot we brute force search over exponentially many subnetworks, testing each on the dataset, and claiming to have found the lottery ticket without training a dense model by the end of this process? The answer to this is that such a search creates high mutual information between the mask and the data, since it involves testing every masked network on the data and choosing the best performing mask. Thus the mutual information based bounds we present in this work hold, because the effective parameter count of the mask derived from this brute force search would be high. Of course, training the full model is a faster way to find a lottery ticket than a brute force search, since training neural networks takes subexponential time (Livni et al., 2014). Algorithms attempting to learn the mask at initialization (Sreenivasan et al., 2022; Ramanujan et al., 2020), before learning weights, are interesting in that this naturally increases $I(\mathbf{m}^{W};\mathcal{D})$ early on in training, although their results are not demonstrated at ImageNet scale, where lots of lottery ticket type results are known to break down (Gale et al., 2019; Frankle et al., 2019).
@@
Necessary conditions. Our results present necessary but not sufficient conditions for smooth interpolation with a parameterized function class. (Bombari et al., 2023) examines whether the overparameterization condition $p=\Omega(nd)$ is sufficient to guarantee smooth interpolation in various neural network regimes, finding that it is not for a random feature network, but is for a network in the NTK regime of neural network training (Jacot et al., 2018).444(Bombari et al., 2023) do not include regularization in their random feature model, whereas (Simon et al., 2023) assume optimal regularization, explaining the seemingly different conclusions about overparameterized RF regression models.
@@
Subtleties Around Effective Parameters. While our idea of effective parameters suggest that one can trade off $I(\mathbf{m}^{W};\mathcal{D})$ and parameter count, this does not imply that a learning algorithm that artificially inflates this mutual information, for instance by encoding a discretization of the data within the mask, will do well as a pruning algorithm. Our modified Law of Robustness presents a lower bound on Lipschitz constant, so that the actual Lipschitz constant could easily be higher than the lower bound given. Additionally, one can view the choice of a mask $\mathbf{m}^{W}$ followed by a choice of weights $\mathbf{w}^{W}$ constrained by this mask as having access to a collection of function classes $\{\mathcal{H}_{t}\}_{t\in T}$, where one first makes a choice $W$ (possibly depending on the data $\mathcal{D}$) of function class $\mathcal{H}_{W}$ and then a choice of function within the class. In such cases, this idea of effective parameterization accounting for both the mutual information of the choice of function class and the parameter count within the chosen class extends beyond the Law of Robustness to other settings (see Theorems C.1, C.2 in Appendix C), showing that the principle we introduce of trading off parameters and information is a general one.
@@ 6 Experiments
Figure 1: Top: memorization capacity (train accuracy on noisy data) against sparsity level for different pruning methods. Staying higher on plots is better. Vertical gap between IMP/Magnitude-after pruning reflect additional memorization capacity on this dataset due to mutual information between mask and data. 2-hidden layer MLP on Gaussian data 1(a) and noisy FashionMNIST 1(b), 4-layer convNet on noisy CIFAR-10 1. Bottom: Ability to correlate with dataset noise over training as a proxy for network capacity. 5-layer ReLU MLP in a student-teacher task, with $\sigma^{2}=1$ noisy labels; 1: correlation with noise during IMP; training increases correlation with noise (hence $p_{\text{eff}}$), pruning then reduces this, before repeating; 1: Sweep over learning rates; 1: Sweep over amount of noise injected into gradients. This illustrates that our quantity of interest, $I(f^{W};\mathcal{D})$, and thus $p_{\text{eff}}$, is increasing due to data contained in gradients.
Motivating the Experimental Setup. The main message of our theory is that pruning at initialization may be inherently difficult because a mask derived through training a dense network contributes to effective parameter count. How can one test whether this occurs in neural networks in practice? Ideally, one would track the mutual information of interest, $I(\mathbf{m^{W}};\mathcal{D})$ throughout training, but because computing mutual information requires estimating high dimensional distributions, it is only exactly feasible in very small models (Kraskov et al., 2004; Gao et al., 2018), which is precisely the setting where pruning to high sparsities leads to layer collapse (Tanaka et al., 2020), where all the weights in one layer vanish. We perform such experiments for a tiny neural network in Appendix B.2. The question then becomes how else one quantify the “effective parameter count” of a network. We resort to tracking two proxies for mutual information and effective parameter count: memorizing capacity and correlation with noise. The first asserts that “effective parameter count” can be measured by tracking “memorization capacity” (Cover, 1965; Sur & Candès, 2019; Montanari et al., 2024). And if one wants to understand “ability to memorize,” using train accuracy noisy data is a natural candidate. Our second proxy for mutual information is more direct: the correlation of the network function with the noise in labels over the course of training. By Lemma 5.2, we have that this correlation is a lower bound for the quantity of interest $I(f^{W};\mathcal{D})$.
To reiterate: what our theory seeks to explain is the empirical finding in (Frankle et al., 2020b) that shows that methods that derive a data-dependent mask (lottery tickets) outperform methods that do not in terms of end-time test accuracy. So any quantity one argues causes this trend in end-time test accuracy (in our case, mutual information and thus effective parameter count) must exhibit the same trend. Our theory predicts mutual information and thus effective parameter count of the sparse network after pruning should be highest for IMP and magnitude-after pruning, and lower for methods that prune at initialization, like SNIP, GraSP, and SynFlow. Our experiments are meant to illustrate that information gained during training can affect model capacity, rather than serve as exhaustive examinations of large modern architectures.
@@ 6.1 Memorization capacity
We consider a two-hidden layer network with ReLU activation, on a train set of points (Gaussian data in Figure 1(a) and FashionMNIST in Figure 1(b)) with noise (random Boolean labels in Figure 1(a), and corrupted images with true labels in Figure 1(b)), as well as a convNet on noisy CIFAR-10 in 1, and compute the fraction of the training set the sparse network can memorize at each sparsity level.
One sees that for all datasets here IMP is more expressive and can memorize more points with the same number of unmasked parameters as the other subnetworks derived from different pruning algorithms. This gap in memorization capacity is due to mutual information with the data, providing evidence that the effective parameter count our theorems reason can indeed correspond to capacity of neural networks in practice. We emphasize that the y-axis is not test accuracy, but memorization capacity (train accuracy on highly noisy data): the fact that our plots give analogous behavior to plots of noiseless test-accuracy against sparsity in, e.g., (Frankle et al., 2020b), is encouraging because the shape of such plots is what we sought to explain. Experimental details are in Appendix B.1.
@@ 6.2 Correlation with noise
We now track our second proxy for mutual information/effective parameter count during training, correlation with label noise. Overall, we see that as the network reaches low train error, it necessarily fits some of the noise, and so correlation with noise increases. We measure this quantity over training in a student-teacher task. Figure 1 illustrates how increasing learning rate increases the rate at which MI is acquired, suggesting that gradients carry this informational quantity. To verify this, we can corrupt the gradients with noise, finding in Figure 1 that this correlational proxy for MI vanishes when we do so. Finally, Figure 1 tracks this proxy for MI during IMP, finding that the MI increases during iterations of pruning, decreases after pruning, and breaks down at late time. In context of our theory where $I(\mathbf{m^{W}};\mathcal{D})$ contributes to $p_{\text{eff}}$, this shows the role of iterative, rather than one-shot pruning, may be to increase the effective parameter count via gradient updates, so it can be pruned back down again. Hence the iterative nature of pruning allows the trade-off between parameters and information to be made many times.
Takeaways. Our experiments illustrate how mutual information can be traded off with parameters so that subnetworks derived from pruning at initialization have a lower effective parameter count than those derived from pruning after training. Our work explains why widespread empirical efforts to prune at initialization have run into difficulties, and introduces a new information-parameter trade-off in doing so that may be of independent interest in understanding neural networks trained by gradient descent.
Impact Statement
Our theoretical results are suggestive about the difficulty of pruning neural networks at initialization. We find that there may be information-theoretic barriers to finding extremely sparse, trainable subnetworks without training the full neural network on data so that research efforts to prune to high sparsities may instead be better directed at pruning jointly with training or before inference, where methods are known to exist for pruning to high sparsities without compromising accuracy. This contributes to a wider literature on the limitations of neural networks.
Appendix A Proof of Theorem 5.6
@@
A.1 Overview of Finite setting, Theorem 5.3
A.1.1 Proof of Lemma 5.4
Lemma A.1.
@@
$A_{f}=\left\{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(x_{i}))^{2}\leq\sigma^{2}-\epsilon\right\}\qquad B$ =$=\left\{\frac{1}{n}\sum_{i=1}^{n}z_{i}^{2}\geq\sigma^{2}-\frac{\epsilon}{6}\right\}\qquad C=\left\{\frac{1}{n}\sum_{i=1}^{n}z_{i}g(x_{i})\geq-\frac{\epsilon}{6}\right\}$
$D_{f}$ =$=\left\{\frac{1}{n}\sum_{i=1}^{n}f(x_{i})z_{i}\geq\frac{\epsilon}{4}\right\}.$
@@
Follows from Lemma 2.1 of (Bubeck & Sellke, 2023). ∎
Lemma A.2.
@@
Follows from Theorem 2 of (Bubeck & Sellke, 2023). ∎
@@
In particular $A_{f^{W}}\implies(B^{c}\cup C^{c})\cup E_{f^{W}}\cup F$, and using the bounds in Lemmas A.1 and A.2 now yields Lemma 5.4.
A.1.2 Proof of Lemma 5.2
Lemma 5.2 follows from Theorem 3 of (Xu & Raginsky, 2017). After translating notation, we have that $\mathbb{P}(|X_{w}|>\alpha)\leq\delta$ whenever
@@
upon which rearranging yields exactly Lemma 5.2.555The explicit constant $a_{1}$ here is not the same as in (Xu & Raginsky, 2017) as a result of different definitions of “subgaussian” - in (Xu & Raginsky, 2017), the authors define $\sigma$-subgaussian to mean a random variable $U$ satisfies $\log\mathbb{E}[e^{\lambda(U-\mathbb{E}U)}]\leq\lambda^{2}\sigma^{2}/2$. The assumed condition $\mathbb{P}(|X|\geq t)\leq 2e^{-t^{2}/C}$, implies $\log\mathbb{E}[e^{\lambda X}]\leq\lambda^{2}(9C)/(2)$ (exercises 3.1 d and e of (van Handel, 2016)) so $a_{1}^{2}=72$.
A.1.3 Isoperimetry, Subgaussianity, and Crude Bounds
Having established Lemma 5.4, to show Theorem 5.3 it remains to upper bound $\mathbb{P}[E_{f^{W}}]$. We first show that $X_{t}=\frac{1}{n}\sum_{i=1}^{n}(f^{t}(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f^{t}])z_{i}$ is subgaussian. As in (Bubeck & Sellke, 2023), the isoperimetry assumption implies that $\sqrt{\frac{d}{c}}\frac{f(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f]}{L}$ is $2$-subgaussian. Since $|z_{i}|\leq 2$, it follows that $\sqrt{\frac{d}{c}}\frac{(f(x_{i})-\mathbb{E}^{\mu_{\ell_{i}}}[f])z_{i}}{L}$ is $8$-subgaussian, by Proposition 1.2 of (Bubeck & Sellke, 2023). Hence each $X_{t}$ is $\left(18\cdot 8\frac{cL^{2}}{dn}\right)$-subgaussian. For convenience we thus define $C_{0}=\frac{144cL^{2}}{dn}$.